OpenStudy (patrickweewoos):

Which equations could model this problem? A 5 lb bag of potatoes costs $3. What is the unit price p per pound? Choose all answers that are correct.

OpenStudy (patrickweewoos):

A.)3/5=p B.)3p=5 C.)5/3=p D.)5p=3

OpenStudy (patrickweewoos):

@iGreen and @perl

OpenStudy (igreen):

To solve you have to divide 5 / 3..so what's your answer? @PatrickWeeWoos

OpenStudy (patrickweewoos):

0.6 and its not tht i typed the answers to the question up there

OpenStudy (patrickweewoos):

Theres multiple ones

OpenStudy (patrickweewoos):

But the answer is 0.6 @iGreen

OpenStudy (igreen):

5 / 3 is not 0.6..

OpenStudy (patrickweewoos):

5 divided by 3 ?

OpenStudy (igreen):

Yes.

OpenStudy (patrickweewoos):

oh its 1.66666667

OpenStudy (igreen):

Yes, but we're not looking for that..which option has 5 / 3 in it? @PatrickWeeWoos

OpenStudy (patrickweewoos):

C.) but its multiple answer question

OpenStudy (igreen):

I think it's B too, because: \(3p = 5\) Divide 3 to both sides: \(p = 5/3\)

OpenStudy (igreen):

So it's B and C.

OpenStudy (patrickweewoos):

Thank you